# Spectral Hardening Reveals Afterglow Emergence in Long-Duration Fast X-ray Transients: A Case Study of GRB 250404A/EP250404a

### v1

@@ 3.1 Early-Time Long-Duration Light Curve
In Figure 1, we present the light curve of GRB 250404A/EP250404a detected by Fermi/GBM with a bin size of 0.5 s in the energy range of 10–1000 keV. The accumulated counts curve is also shown, from which we extract the $T_{90}$ interval of ${90.43}_{-0.37}^{+0.64}$ s. The prompt emission phase of GRB 250404A is indicated by the $T_{90}$ and multi-pulse light curve with variability in this energy range.
@@
We also display the scaled long-duration light curve detected by EP/WXT with a bin size of 3 s in the energy range of 0.5–4.0 keV. Continuous X-ray emission was affected by the slewing motion from $T_{0}$ + 15 s to 130 s and was disrupted at $T_{0}$ + 1384 s due to the end of the observation cycle. As a result, a reliable $T_{90}$ estimation in X-ray band for EP250404a could not be accurately derived. We also note that the second pulse of EP250404a appears broadened and extended compared to the gamma-ray emission, exhibiting a long tail with continuous emission that persists for thousands of seconds.
@@
3.2 Spectral Evolution
We performed both time-integrated and time-resolved spectral fittings for Fermi/GBM and EP data using the Python package bayspec [^2] https://github.com/jyangch/bayspec (an upgraded version of MySpecFit), following the approach described in Yang et al. (2022) and Yang et al. (2023). bayspec is a Bayesian inference-based spectral fitting tool for multi-dimensional and multi-wavelength astrophysical data. The goodness of fit was evaluated by examining the reduced statistic STAT/dof, as described in Yin et al. (2024). The model comparison was conducted using the Bayesian information criterion (BIC) as defined by Schwarz (1978). The best-fitting model parameters, along with the corresponding statistics for each time slice, are provided in Table 3 and Table 4. Based on the spectral fitting results, we derived the spectral evolution and the spectral energy distributions (SEDs), which are illustrated in Figure 1.
Before analyzing the spectral evolution in detail, we firstly divided the early high-energy emissions detected by Fermi/GBM and EP into four time intervals based on data availability and performed time-integrated spectral fittings for each episode:
  • (a)
    EP/WXT and Fermi/GBM joint fit in $T_{0}$ + [-1, 15] s. Before the time interval when EP was slewing, there was a single overlapping time range of gamma-ray and X-ray emission, spanning $T_{0}$ + [-1, 15] s. We performed a joint fit using an absorbed cutoff power law (CPL) model, tbabs*ztbabs*cpl. In this model, tbabs and ztbabs represent the Tuebingen-Boulder ISM absorption model (Wilms et al., 2000), with parameters for the absorption column density, $N_{\rm H}$, and redshift. We adopted a Galactic absorption column density of $N_{\rm H,gal}\sim 6.00\times 10^{20}\leavevmode\nobreak\ \rm{cm^{-2}}$ and an intrinsic absorption column density of $N_{\rm H,int1}\sim 3.73\times 10^{22}\leavevmode\nobreak\ \rm{cm^{-2}}$, with the latter fixed to the value obtained from the time-integrated spectral fitting in $T_{0}$ + [130, 255] s using the EP/FXT spectrum (see below). The joint spectrum is well fitted, yielding a spectral index of ${-0.93}_{-0.04}^{+0.05}$ and a peak energy of $50.93_{-1.62}^{+1.55}$ keV. This time interval covers the first pulse of the prompt emission detected in gamma-rays, and the joint best-fit parameters represent an averaged view of the time-resolved spectral evolution during the first pulse (see below).
  • (b)
    Fermi/GBM independent fit in $T_{0}$ + [-4, 113] s. The prompt emission of GRB 250404A detected by Fermi/GBM expands from $\sim$ $T_{0}$4 s to +113 s. The time-integrated spectrum for this interval is well described by the CPL model, yielding a spectral index of ${-1.15}_{-0.02}^{+0.02}$ and a peak energy of ${55.34}_{-0.51}^{+0.64}$ keV. The isotropic energy derived from this episode is $3.05_{-0.01}^{+0.01}\times 10^{53}$ erg, placing GRB 250404A within the parameter space of Type II GRBs on the Amati relation (Amati et al., 2002).
  • (c)
    EP/FXT independent fit in $T_{0}$ + [130, 255] s and [255, 1384] s. The long-duration tail of EP250404a was detected by EP/FXT from $T_{0}$ + 130 s to + 1384 s. Considering the possibility of time-dependent absorption in the X-ray spectra, which is not uncommon in bright GRBs (Lazzati & Perna, 2002; Perna et al., 2003; Lazzati & Perna, 2003; Campana et al., 2007, 2021), we adopted different intrinsic absorption column densities for the time-resolved spectral fittings. However, the values of $N_{\rm H,int}$ for the time-resolved spectra are difficult to constrain accurately, with significant uncertainties. Thus, we determined the $N_{\rm H,int}$ from time-integrated spectra and fixed the value for time-resolved spectral fittings.
    We firstly performed a time-resolved spectral fitting test using the model tbabs*ztbabs*pl. In this model, the Galactic absorption column density $N_{\rm H,gal}$ was fixed at $6.00\times 10^{20}\leavevmode\nobreak\ \rm{cm^{-2}}$, while the intrinsic absorption column density $N_{\rm H,int}$ was treated as a free parameter. We divided the time-resolved spectra, ensuring each EP/FXT spectrum contained at least 300 total accumulating photon counts. Our analysis revealed significant evolution in the best-fit values of $N_{\rm H,int}$ between the fast-decaying phase ($\sim T_{0}$ + [130, 255] s) and the slow-decaying phase ($\sim T_{0}$ + [255, 1384] s) of the X-ray emission tail. Therefore, we adopted two distinct intrinsic absorption column densities determined from independent fit in $T_{0}$ + [130, 255] s and [255, 1384] s: $N_{\rm H,int1}$ $\sim$ $3.73\times 10^{22}$ $\rm cm^{-2}$ for the period before $T_{0}$ + 255 s and $N_{\rm H,int2}$ $\sim$ $6.39\times 10^{21}$ $\rm cm^{-2}$ for the period afterward.
    Accounting for the effect of an evolving $N_{\rm H,int}$ is non-trivial because the absorption in X-ray spectra is degenerate with the intrinsic spectral index. As shown in Figure 2, our treatment of the intrinsic absorption column density does not significantly affect the values of the spectral index, nor does it alter the overall trend of spectral index evolution.
Building on the time-integrated spectral fitting results, we performed time-resolved spectral fittings to derive the spectral evolution and SEDs, as shown in Figure 1. We briefly outline the time-resolved spectral fittings and analyses as follows:
@@ 3.2 Spectral Evolution
    The prompt emission in gamma-rays.
    We extracted the time-resolved spectra of Fermi/GBM in the time range of $T_{0}$ + [-4, 113] s to explore the spectral evolution of the prompt emission in gamma-rays. The time-resolved spectra were divided into two sets based on different criteria: one with a minimum of 30 average accumulated photon counts per channel and the other with a minimum of 120. The extracted spectra were fitted using the CPL model, and the spectral evolution was analyzed through the best-fit parameter. The resulting spectral evolution and SED plots are shown in Figure 1, where we observe that the peak energy, $E_{\rm p}$, exhibits a hard-to-soft transition during the first pulse, followed by an intensity tracking pattern (Golenetskii et al., 1983), with an overall declining trend. Additionally, the plots reveal that the spectral index, $\alpha_{\gamma}$, evolves with an overall softening trend, ranging from -0.51 to -1.58.
@@
    The spectral evolution of the long-duration tail in X-rays was analyzed via time-resolved spectral fittings with fixed intrinsic absorption column density $N_{\rm H,int1}$ and $N_{\rm H,int2}$ determined from time-integrated spectral fittings. We employed two spectral models, an absorbed powerlaw (PL) tbabs*ztbabs*pl and an absorbed smooth broken powerlaw (SBPL) tbabs*ztbabs*sbpl, for the time-resolved spectral fittings of EP/FXT spectra. Most spectra were better described by tbabs*ztbabs*pl. However, for the time-resolved spectra in the intervals $T_{0}$ + [130, 154] s and $T_{0}$ + [290.5, 462.0] s, both models provided comparable fits ($\Delta\rm{BIC}\textless 5$). We identified the spectral evolution of the peak energy passing through EP/FXT 0.5–10.0 keV band within $T_{0}$ + [130, 154] s, as revealed by the SBPL model, which aligns with the softening trend observed in Fermi/GBM spectral fittings in $T_{0}$ + [-4, 113] s. Furthermore, the spectral indices $\alpha_{\rm X}$ demonstrate significant hardening after $T_{0}$ + 255 s, eventually reaching a plateau at $\sim-2$. Notably, the first two time-resolved spectra after $T_{0}$ + 255 s clearly display a transition, as seen in the SBPL model in the SED plot. This suggests the emergence of a second spectral component, distinct from the softening and fading prompt emission.
@@
Lastly, joint fittings were performed using EP/FXT-A and EP/FXT-B spectra for the time intervals $T_{0}$ + [44412, 47458] s and $T_{0}$ + [114594, 122326] s. We incorporated a calibration constant for EP/WXT-A to account for the systematic uncertainty between the two detectors. The X-ray flux densities were then derived for the afterglow fitting.
Figure 3: Multiwavelength observations of GRB 250404A/EP250404a and afterglow modeling with best-fit parameters from the FS+RS model with $u$, $v$, and $g$ band correction factors. The multiwavelength data utilized in the afterglow fitting are marked in solid points, while the optical observations on clear filters that are not included in the afterglow fitting are marked in hollow points. The optical and near-infrared data have been corrected for Galactic extinction, which is $E(B-V)=0.073$ mag (Schlafly & Finkbeiner, 2011). We assume a total to selective extinction ratio of $R_{V}=3.1$ according to the extinction law from Fitzpatrick (1999). The inverted triangle points signify upper limits. The best-fit FS+RS, FS and RS models are shown with solid, dashed and dotted lines, respectively. The dark blue dashed vertical lines corresponds to the same time interval where the spectra start to harden and reach a plateau.
Figure 4: X-ray/optical/near-infrared afterglow specific flux density spectra of GRB 250404A/EP250404a in different time intervals. The best-fit afterglow model is indicated by dashed gray lines. The red curves correspond to the partial host galaxy extinction curve derived from the best-fit correction factors from the afterglow modeling. For comparison, the yellow, blue and green curves rshow the average extinction laws of the MW, SMC, LMC, respectively, assuming a reddening value of $E(B-V)=0.1$ mag.
3.3 Confirming Afterglow Emergence via Multiwavelength Fitting
Corresponding to the spectral hardening phase starting at $\sim$ $T_{0}$ + 255 s, the flux density at 1 keV initially exhibits a brief decay, followed by a sharp rise peaking at $\sim$ $T_{0}$ + 1000 s. This rising flux is also observed in the $u$, $g$ and $i$ band of Mephisto early-time optical data. Subsequently, the flux density shows a steeper slope before $\sim$ $T_{0}$ + 5000 s and a shallower decay at later times. These behaviors indicate the presence of a RS component emerging around $T_{0}$ + 1000 s, superimposed on the FS emission.
We fitted the multiwavelength data with both FS and FS+RS models using PyFRS [^3] https://github.com/leiwh/PyFRS (Zhang, 2018; Gao et al., 2013; Lei et al., 2016; Zhu et al., 2023). Given the potential host galaxy extinction affecting the blue bands, we adopted two approaches: (1) fitting the afterglow without the heavily affected $u$ and $v$ bands, and (2) fitting the afterglow with extinction correction factors applied to the $u$, $v$, and $g$ bands.
The fitting of the multiwavelength data is implemented by the Bayesian computation python package PyMultinest (Buchner et al., 2014) with the log-likelihood function written as
where $O_{i}$, $P_{i}$, and $\sigma_{i}$ stand for the $i$th of $n$ observed magnitudes, model-predicted magnitudes, and the uncertainties of observed magnitudes, respectively. An additional variance parameter $v$ is introduced to as a scatter term, which accounts for additional uncertainty in models and data. For upper limits, a one-sided Gaussian penalty term is applied. To measure the goodness of fit between the model and the observed data, we utilize $\chi^{2}$ as the statistical metric. All model parameters are allowed to vary, with broad but physically reasonable priors, enabling a comprehensive search for the best-fit solution.
The best-fitting model parameters and their corresponding statistics for the four fitting scenarios are presented in Table 2. Among these, only the FS+RS model with extinction correction factors accurately reproduces the multiwavelength data. The corner plot of the posterior probability distributions of the parameters are presented in Figure 7. Furthermore, the derived extinction correction factors are consistent with the specific flux density spectra shown in Figure 4 in different time slices, which formed a steeper curve compared to the average extinction laws of Milky Way (MW), Small Magellanic Clouds (SMC) and Large Magellanic Clouds (LMC), assuming a reddening value of $E(B-V)=0.1$.

### v2

@@ 3.1 Early-time Long-duration Light Curve
In Figure 1, we present the light curve of GRB 250404A/EP250404a detected by Fermi/GBM with a bin size of 0.5 s in the energy range of 10–1000 keV. The accumulated counts curve is also shown, from which we extract the $T_{90}$ interval of ${90.43}_{-0.37}^{+0.64}$ s. The prompt emission phase of GRB 250404A is indicated by the $T_{90}$ and multipulse light curve with variability in this energy range.
@@
We also display the scaled long-duration light curve detected by EP/WXT with a bin size of 3 s in the energy range of 0.5–4.0 keV. Continuous X-ray emission was affected by the slewing motion from $T_{0}$ + 15 s to 130 s and was disrupted at $T_{0}$ + 1384 s due to the end of the observation cycle. As a result, a reliable $T_{90}$ estimation in the X-ray band for EP250404a could not be accurately derived. We also note that the second pulse of EP250404a appears broadened and extended compared to the gamma-ray emission, exhibiting a long tail with continuous emission that persists for thousands of seconds.
3.2 Spectral Evolution
We performed both time-integrated and time-resolved spectral fittings for Fermi/GBM and EP data using the Python package bayspec [^2] https://github.com/jyangch/bayspec (an upgraded version of MySpecFit), following the approach described in J. Yang et al. (2022) and J. Yang et al. (2023). bayspec is a Bayesian-inference-based spectral fitting tool for multidimensional and multiwavelength astrophysical data. The goodness of fit was evaluated by examining the reduced statistic STAT/dof, as described in Y.-H. I. Yin et al. (2024). The model comparison was conducted using the Bayesian information criterion (BIC) as defined by G. Schwarz (1978). The best-fitting model parameters, along with the corresponding statistics for each time slice, are provided in Tables 3 and 4. Based on the spectral fitting results, we derived the spectral evolution and the spectral energy distributions (SEDs), which are illustrated in Figure 1.
Before analyzing the spectral evolution in detail, we firstly divided the early high-energy emissions detected by Fermi/GBM and EP into four time intervals based on data availability and performed time-integrated spectral fittings for each episode.
  • (a)
    EP/WXT and Fermi/GBM joint fit in $T_{0}$ + [-1, 15] s. Before the time interval when EP was slewing, there was a single overlapping time range of gamma-ray and X-ray emission, spanning $T_{0}$ + [-1, 15] s. We performed a joint fit using an absorbed cutoff power-law (CPL) model, tbabs*ztbabs*cpl. In this model, tbabs and ztbabs represent the Tuebingen-Boulder interstellar medium (ISM) absorption model (J. Wilms et al., 2000), with parameters for the absorption column density, $N_{\rm H}$, and redshift. We adopted a Galactic absorption column density of $N_{\rm H,gal}\sim 6.00\times 10^{20}\leavevmode\nobreak\ \rm{cm^{-2}}$ and an intrinsic absorption column density of $N_{\rm H,int1}\sim 3.73\times 10^{22}\leavevmode\nobreak\ \rm{cm^{-2}}$, with the latter fixed to the value obtained from the time-integrated spectral fitting in $T_{0}$ + [130, 255] s using the EP/FXT spectrum (see below). The joint spectrum is well fitted, yielding a spectral index of ${-0.93}_{-0.04}^{+0.05}$ and a peak energy of $50.93_{-1.62}^{+1.55}$ keV. This time interval covers the first pulse of the prompt emission detected in gamma rays, and the joint best-fit parameters represent an averaged view of the time-resolved spectral evolution during the first pulse (see below).
  • (b)
    Fermi/GBM independent fit in $T_{0}$ + [-4, 113] s. The prompt emission of GRB 250404A detected by Fermi/GBM expands from $\sim$ $T_{0}$ -4 s to +113 s. The time-integrated spectrum for this interval is well described by the CPL model, yielding a spectral index of ${-1.15}_{-0.02}^{+0.02}$ and a peak energy of ${55.34}_{-0.51}^{+0.64}$ keV. The isotropic energy derived from this episode is $3.05_{-0.01}^{+0.01}\times 10^{53}$ erg, placing GRB 250404A within the parameter space of type II GRBs on the Amati relation (L. Amati et al., 2002).
  • (c)
    EP/FXT independent fit in $T_{0}$ + [130, 255] s and [255, 1384] s. The long-duration tail of EP250404a was detected by EP/FXT from $T_{0}$ + 130 s to + 1384 s. Considering the possibility of time-dependent absorption in the X-ray spectra, which is not uncommon in bright GRBs (D. Lazzati & R. Perna, 2002; R. Perna et al., 2003; D. Lazzati & R. Perna, 2003; S. Campana et al., 2007, 2021), we adopted different intrinsic absorption column densities for the time-resolved spectral fittings. However, the values of $N_{\rm H,int}$ for the time-resolved spectra are difficult to constrain accurately, with significant uncertainties. Thus, we determined the $N_{\rm H,int}$ from time-integrated spectra and fixed the value for time-resolved spectral fittings.
    We firstly performed a time-resolved spectral fitting test using the model tbabs*ztbabs*pl. In this model, the Galactic absorption column density $N_{\rm H,gal}$ was fixed at $6.00\times 10^{20}\leavevmode\nobreak\ \rm{cm^{-2}}$, while the intrinsic absorption column density $N_{\rm H,int}$ was treated as a free parameter. We divided the time-resolved spectra, ensuring each EP/FXT spectrum contained at least 300 total accumulating photon counts. Our analysis revealed significant evolution in the best-fit values of $N_{\rm H,int}$ between the fast-decaying phase ($\sim T_{0}$ + [130, 255] s) and the slow-decaying phase ($\sim T_{0}$ + [255, 1384] s) of the X-ray emission tail. Therefore, we adopted two distinct intrinsic absorption column densities determined from independent fits in $T_{0}$ + [130, 255] s and [255, 1384] s: $N_{\rm H,int1}$ $\sim$ $3.73\times 10^{22}$ $\rm cm^{-2}$ for the period before $T_{0}$ + 255 s and $N_{\rm H,int2}$ $\sim$ $6.39\times 10^{21}$ $\rm cm^{-2}$ for the period afterward.
    Accounting for the effect of an evolving $N_{\rm H,int}$ is nontrivial because the absorption in X-ray spectra is degenerate with the intrinsic spectral index. As shown in Figure 4, our treatment of the intrinsic absorption column density does not significantly affect the values of the spectral index, nor does it alter the overall trend of spectral index evolution.
@@ 3.2 Spectral Evolution
Building on the time-integrated spectral fitting results, we performed time-resolved spectral fittings to derive the spectral evolution and SEDs, as shown in Figure 1. We briefly outline the time-resolved spectral fittings and analyses as follows.
@@
    The prompt emission in gamma rays.
    We extracted the time-resolved spectra of Fermi/GBM in the time range of $T_{0}$ + [-4, 113] s to explore the spectral evolution of the prompt emission in gamma rays. The time-resolved spectra were divided into two sets based on different criteria: one with a minimum of 30 average accumulated photon counts per channel and the other with a minimum of 120. The extracted spectra were fitted using the CPL model, and the spectral evolution was analyzed through the best-fit parameter. The resulting spectral evolution and SED plots are shown in Figure 1, where we observe that the peak energy, $E_{\rm p}$, exhibits a hard-to-soft transition during the first pulse, followed by an intensity tracking pattern (S. V. Golenetskii et al., 1983), with an overall declining trend. Additionally, the plots reveal that the spectral index, $\alpha_{\gamma}$, evolves with an overall softening trend, ranging from -0.51 to -1.58.
@@
    The spectral evolution of the long-duration tail in X-rays was analyzed via time-resolved spectral fittings with fixed intrinsic absorption column density $N_{\rm H,int1}$ and $N_{\rm H,int2}$ determined from time-integrated spectral fittings. We employed two spectral models, an absorbed power law (PL) tbabs*ztbabs*pl and an absorbed smooth broken power law (SBPL) tbabs*ztbabs*sbpl, for the time-resolved spectral fittings of EP/FXT spectra. Most spectra were better described by tbabs*ztbabs*pl. However, for the time-resolved spectra in the intervals $T_{0}$ + [130, 154] s and $T_{0}$ + [290.5, 462.0] s, both models provided comparable fits ($\Delta\rm{BIC}\textless 5$). We identified the spectral evolution of the peak energy passing through the EP/FXT 0.5–10.0 keV band within $T_{0}$ + [130, 154] s, as revealed by the SBPL model, which aligns with the softening trend observed in Fermi/GBM spectral fittings in $T_{0}$ + [-4, 113] s. Furthermore, the spectral indices $\alpha_{\rm X}$ demonstrate significant hardening after $T_{0}$ + 255 s, eventually reaching a plateau at $\sim-2$. Notably, the first two time-resolved spectra after $T_{0}$ + 255 s clearly display a transition, as seen in the SBPL model in the SED plot. This suggests the emergence of a second spectral component, distinct from the softening and fading prompt emission.
@@
Lastly, joint fittings were performed using EP/FXT-A and EP/FXT-B spectra for the time intervals $T_{0}$ + [44412, 47458] s and $T_{0}$ + [114594, 122326] s. We incorporated a calibration constant for EP/FXT-A to account for the systematic uncertainty between the two detectors. The X-ray flux densities were then derived for the afterglow fitting.
@@
3.3 Confirming Afterglow Emergence via Multiwavelength Fitting
Corresponding to the spectral hardening phase starting at $\sim$ $T_{0}$ + 255 s, the flux density at 1 keV initially exhibits a brief decay, followed by a sharp rise peaking at $\sim$ $T_{0}$ + 1000 s. This rising flux is also observed in the $u$, $g$ and $i$ bands of Mephisto early-time optical data. Subsequently, the flux density shows a steeper slope before $\sim$ $T_{0}$ + 5000 s and a shallower decay at later times. These behaviors indicate the presence of a RS component emerging around $T_{0}$ + 1000 s, superimposed on the FS emission.
We fitted the multiwavelength data with both FS and FS+RS models using PyFRS [^3] https://github.com/leiwh/PyFRS (B. Zhang, 2018; H. Gao et al., 2013; W.-H. Lei et al., 2016; Z.-P. Zhu et al., 2023). In our modeling, we adopt the top-hat jet-type and thin-shell case. The overall evolution of the FS shell, covering the coasting, deceleration, jet-break, and Newtonian phases, is calculated numerically using a set of hydrodynamical equations as given in Y. F. Huang et al. (2000). In the thin shell case, the RS is Newtonian during the shock crossing phase, and its dynamical behavior is expressed with the scalings as described in S. Kobayashi (2000) (see also H. Gao et al. (2013); B. Zhang (2018)). [^4] It should be noted that the dynamics in PyFRS are approximate descriptions. A more rigorous treatment accounting for internal energy evolution, adiabatic losses, and total energy conservation is given by L. Nava et al. (2013) (see also B. Zhang (2018)) The synchrotron spectra of the FS and RS are calculated following the standard broken power-law spectral model, separated by the three characteristic frequencies: the synchrotron self-absorption frequency $\nu_{a}$, the minimum frequency $\nu_{m}$, and the cooling frequency $\nu_{c}$. In parallel with the physical modeling, we also considered the potential host galaxy extinction affecting the blue bands. Here, we adopted two approaches: (1) fitting the afterglow without the heavily affected $u$ and $v$ bands and (2) fitting the afterglow with extinction correction factors applied to the $u$, $v$, and $g$ bands.
The fitting of the multiwavelength data is implemented by the Bayesian computation python package PyMultinest (J. Buchner et al., 2014) with the log-likelihood function written as
where $O_{i}$, $P_{i}$, and $\sigma_{i}$ stand for the $i$th of $n$ observed magnitudes, model-predicted magnitudes, and the uncertainties of observed magnitudes, respectively. An additional variance parameter $v$ is introduced as a scatter term, which accounts for additional uncertainty in the models and data. For upper limits, a one-sided Gaussian penalty term is applied. To measure the goodness of fit between the model and the observed data, we utilize $\chi^{2}$ as the statistical metric. All model parameters are allowed to vary, with broad but physically reasonable priors, enabling a comprehensive search for the best-fit solution.
The best-fitting model parameters and their corresponding statistics for the four fitting scenarios are presented in Table 2. Among these, only the FS+RS model with extinction correction factors accurately reproduces the multiwavelength data. The corner plot of the posterior probability distributions of the parameters are presented in Figure 7. Furthermore, the derived extinction correction factors are consistent with the specific flux density spectra shown in Figure 3 in different time slices, which formed a steeper curve compared to the average extinction laws of Milky Way (MW), Small Magellanic Cloud (SMC) and Large Magellanic Cloud (LMC), assuming a reddening value of E(BV)=0.1E(B-V)=0.1.